OpenStudy (anonymous):

2(2r-1)+4=5(r+1) solve the linear equation

OpenStudy (lgbasallote):

distribute the distributable first.. \[\implies 4r - 2 + 4 = 5r + 5\] combine like terms... \[\implies 4r + 2 = 5r + 5\] subtract 4r from both sides \[\implies 2 = 5r - 4r + 5\] subtract 5 from both sides \[\implies 2 - 5 = 5r - 4r\] does that help?